Circular No. 6150

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

GRO J1849–03 S. N. Zhang, B. A. Harmon, M. H. Finger, G. J. Fishman, W. S.

Paciesas, and C. A. Wilson, Marshall Space Flight Center, NASA; J. E. Grindlay, D. Barret, and P. Bloser, Harvard-Smithsonian Center for Astrophysics; and E. Ford, M. Tavani, and P. Kaaret, Columbia University, report: “Analysis of Compton Observatory BATSE data has revealed a possible new x-ray binary system. We have observed six outbursts from the source (GRO J1849–03), spaced at 241-day (± 1) intervals in the period 1991 Apr.–1994 Dec. within an error box α = 18^{h}49^{m} ± 1^{m}, δ = −3^{o}22′ ± 30′ (equinox 2000.0). The epoch-folded light curve of the six periodic outbursts has a FWHM of 13 days and a maximum on 1991 June 5. All six outbursts show similar light curves and spectral behavior. From 20 to 100 keV, the peak flux is 75 ± 20 mCrab, and the spectrum can be fit by a power-law with photon index of −2.8 ± 0.2. We cannot positively identify this source with any of the nearby hard-x-ray sources (Parmar et al. 1993, A.Ap. 279, 179; Koyama et al. 1990, Nature 343, 148) but await a more precise source position that might be obtained with x-ray or possibly optical/infrared observations from the next predicted outburst. An outburst previously reported (IAUC 6096) from this location is consistent with the light curves of the other five outbursts; therefore we now believe that to be associated with this new source rather than EXO 1846–031. We interpret the 241-day period as the orbital period of a binary system. If the periodic outbursts continue, the next outburst should peak on 1995 May 21. Observations are encouraged during the next outburst, especially soft-x-ray and optical/infrared observations for a more accurate location and possible identification of this system.”

α ORIONIS

E. Guinan and D. Steelman, Villanova University, report that Betelgeuse is currently undergoing another major dimming episode. Since 1994 Sept., its brightness at optical wavelengths has decreased by ∼ 0.4 mag. When last measured on 1995 Mar. 10, it was at V = +0.80.

SV SAGITTAE L. Szentaskó, Veresegyház, Hungary, reports that this R CrB-type vari-

able has faded, as indicated by his visual magnitude estimates: 1994 Dec. 31.69 UT, 11.1; 1995 Jan. 2.69, 11.7; Feb. 1.18, 13.2; Mar. 9.09, 13.1.

1995 March 16 Daniel W. E. Green